# Prox-DBRO-VR: A Unified Analysis on Decentralized Byzantine-Resilient Composite Stochastic Optimization with Variance Reduction and Non-Asymptotic Convergence Rates

### v6

@@ IV-B Main Results
(Bounded Gradient). The sub-gradient $\partial_{\tilde{x}}g\left({\tilde{x}}\right)$ at any point ${\tilde{x}}\in\mathbb{R}^{n}$ is $\hat{G}$-bounded, i.e., $\left\|{{\partial_{\tilde{x}}}g\left({\tilde{x}}\right)}\right\|_{2}^{2}\leq\hat{G}$, where $\hat{G}$ can be an arbitrarily large but finite positive constant.
To proceed, we define $\theta>4/\gamma$, $P_{1}^{d}:=16{\left|\mathcal{R}\right|}\hat{G}{\rm{+}}16n{\bar{\phi}^{2}}\sum\nolimits_{i\in{\cal R}}{{{\left|{{{\cal R}_{i}}}\right|}^{2}}}{\rm{+4}}n{\bar{\phi}^{2}}\sum\nolimits_{i\in{\cal R}}{{{\left|{{{\cal B}_{i}}}\right|}^{2}}}$, $\Xi:=\max\left\{{\frac{{{\theta^{2}}P_{1}^{d}}}{{\gamma\theta-1}},\left({\xi-\frac{\gamma}{4}\theta}\right){\left\|{{x_{0}}-{x^{*}}}\right\|_{2}^{2}}+\frac{{{\theta^{2}}}}{\xi}P_{1}^{d}+\theta{P_{2}}-\xi E}\right\}$, and $\xi:={\kappa_{q}}\left({64{{\left({1+{\kappa_{f}}}\right)}^{2}}+{q_{\min}}}\right)\mu\theta$.
@@
The convergence results established in Theorems 2-3 assert that the proposed algorithms achieve linear convergence at the expense of a larger larger convergence error. It is also clear from Theorem 3 that the convergence error of Prox-DBRO-SAGA and Prox-DBRO-LSVRG for the sub-linear convergence case is determined by the number of Byzantine agents. That is to say, the exact convergence of Prox-DBRO-SAGA and Prox-DBRO-LSVRG can be recovered, when the number of Byzantine agents goes to zero. However, the convergence error of Prox-DBRO-SAGA and Prox-DBRO-LSVRG is determined jointly by the number of Byzantine agents and reliable agents in the linear convergence case.
V Experimental Results
In this section, we perform a case study on decentralized soft-max regression with sparsity to verify the theoretical results and show the convergence performance of Prox-DBRO-SAGA and Prox-DBRO-LSVRG, where three kinds of Byzantine attacks (zero-sum attacks, Gaussian attacks, and same-value attacks) are considered. The communication networks are randomly generated by the Erdős-Rényi method, where Byzantine agents are also selected in a random way. All simulations are carried out in Python (version 3.9) on a DELL server (Linux) with 20 Cores 40 Threads i9-10900X 3.70 GHz processor and 32GB memory.
@@ V Experimental Results
In simulation, existing literature adopts only testing accuracy and the consensus error to validate the convergence performance of the proposed algorithms. However, both a higher testing accuracy and a smaller consensus error cannot reflect the convergence of tested algorithms in a complete way. There is a gap between theoretical results and practical applications, especially when facing machine-learning problems. To bridge the gap, we introduce the third performance index: optimality gaps, which respects the theoretical convergence results and could better characterize the transient behavior (convergence or divergence) of algorithms when training a machine-learning model. We consider three networked multi-agent systems under zero-sum attacks, Gaussian attacks, and same-value attacks. Specifically, a network of $m$ agents, consisting of ${\left|\mathcal{R}\right|}$ reliable agents and $\left|\mathcal{B}\right|$ Byzantine agents, minimize a regularized soft-max regression problem for a multi-class classification task via specifying the problem formulation (1) as ${f_{i}}\left({\tilde{x}}\right):=-\left({1/{q_{i}}}\right)\sum\nolimits_{j=1}^{{q_{i}}}{\sum\nolimits_{l=0}^{\tilde{C}-1}{{\mathcal{I}_{\left({{b_{ij}}=l}\right)}}}}\ln\left({{e^{\left[{\tilde{x}}\right]_{l}^{\top}{c_{ij}}}}/\sum\nolimits_{t=0}^{\tilde{C}-1}{{e^{\left[{\tilde{x}}\right]_{t}^{\top}{c_{ij}}}}}}\right)+\left({{\beta_{1}}/2}\right)\left\|{\tilde{x}}\right\|_{2}^{2}$ and ${g}\left({\tilde{x}}\right):={\beta_{2}}{\left\|{\tilde{x}}\right\|_{1}}$, where $\tilde{x}\in{\mathbb{R}^{n}}$ with $n=\tilde{C}\tilde{n}$ is the model parameter, ${\tilde{C}}$ represents the number of classes of total samples, ${\left[{\tilde{x}}\right]_{j}}$ denotes a vector that contains $j\tilde{n}$-th to $\left(\left({j+1}\right)\tilde{n}-1\right)$-th elements of $\tilde{x}$, ${{b_{ij}}}$ and ${{c_{ij}}}$ are the $j$-th label and image allocated to agent $i$, respectively, ${{\mathcal{I}_{\left({{b_{ij}}=l}\right)}}}$ is the indicator function with ${\mathcal{I}_{\left({{b_{ij}}=l}\right)}}=1$ if ${{b_{ij}}=l}$ and ${\mathcal{I}_{\left({{b_{ij}}=l}\right)}}=0$ otherwise, ${\beta_{1}}$ and ${\beta_{2}}$ are positive parameters of the corresponding regularized terms for avoiding over-fitting and obtaining a sparse solution, respectively. The regularized parameters are set as ${\beta_{1}}={\beta_{2}}=1/N$ in following simulations. Since the algorithmic framework BRIDGE [23] is only available to smooth single optimization problems, we equip BRIDGE with the proximal-gradient mapping method to obtain Prox-BRIDGE-T, Prox-BRIDGE-M and Prox-BRIDGE-K for the non-smooth composite optimization problem, which is also applied to GeoMed [51] to get Prox-GeoMed. The initial status of decision variables of all tested algorithms are the same and generated from a standard normal distribution. A total number of $Q=60000$ samples randomly selected from the MNIST [52] dataset are evenly allocated to each agent (including both reliable agents and Byzantine agents over network) to train the discriminator, while the rest of samples are used for testing. Fig. 1 presents 100 samples randomly selected from the dataset.
@@
Zero-sum attacks: As depicted in Fig. (2a), an $m=30$ multi-agent CPS consists of ${\left|\mathcal{R}\right|}=25$ reliable agents (yellow nodes) and $\left|\mathcal{B}\right|=5$ Byzantine agents (red nodes), where each Byzantine agent $o$, $o\in\mathcal{B}$, sends a well-designed malicious message ${z_{io,k}}=-{\sum\nolimits_{j\in{\mathcal{R}_{i}}}{{w_{ij}}{x_{j,k}}}}/\left|{{\mathcal{B}_{i}}}\right|/{w_{oi}}$ to its reliable neighbor $i$, $i\in{\mathcal{R}_{o}}$, to let the states of the reliable agent ${x_{i,k}}=0_{n}$ at each iteration. For NIDS [37], we choose the algorithm parameter $c=1/\left({2{{\max}_{i\in\mathcal{R}}}\left\{{{\alpha_{i}}}\right\}}\right)$ for $\tilde{W}={I_{m}}-c{D_{\alpha}}\left({{I_{m}}-W}\right)$, where $\tilde{W}$, $W$, and $D_{\alpha}$ are the modified mixing matrix, weight matrix, and uncoordinated step-size, respectively. This means if $c$ is sufficiently small, NIDS runs without any communication happening among all agents (both reliable agents and Byzantine agents) over networks. For PMGT-SAGA/PMGT-LSVRG [35], we hand-tune the parameter associated with multi-step communications to obtain the best performance. It is clear from Figs. (2b)-(2d) and Table II that both Prox-DBRO-SAGA and Prox-DBRO-LSVRG achieve a smaller optimality gap and higher testing accuracy than the other tested algorithms in a same amount of computational costs (epochs). This demonstrates that Prox-DBRO-SAGA and Prox-DBRO-LSVRG approximate faster to the optimal solution than the other tested algorithms. It is worthwhile to mention that zero-sum attacks launched by Byzantine agents aim to drive the states of all reliable agents to zero in each iteration. Therefore, the much smaller consensus error of NIDS and PMGT-SAGA/PMGT-LSVRG than the other decentralized Byzantine-resilient algorithms, indicates reversely that they are vulnerable to the zero-sum attacks. Likewise, the bigger consensus error of Prox-DBRO-SAGA and Prox-DBRO-LSVRG means that they are more robust to the zero-sum attacks than the other tested algorithms.
@@
(b) Optimality gaps over epochs.
@@
(d) Consensus errors over epochs.
Figure 2: Performance of all tested algorithms over a Byzantine network under zero-sum attacks.
S&PI Algorithms NIDS PMGT-LSVRG PMGT-SAGA Prox-BRIDGE-T Prox-BRIDGE-M Prox-BRIDGE-K Prox-GeoMed Prox-DBRO-LSVRG Prox-DBRO-SAGA
Step-Size [0.01, 0.015] 0.001 0.01 0.35 0.3 0.35 0.35 0.05 0.005
Triggered Probability N/A $m/Q$ N/A N/A N/A N/A N/A $\left[{m/Q/2,m/Q}\right]$ N/A
Penalty Parameter N/A N/A N/A N/A N/A N/A N/A [0.2, 0.25] [0.2, 0.25]
Consensus Error 0 1.2650e-06 1.2621e-06 0.0012 0.0019 0.0019 0.0010 4.4680 0.0428
Testing Accuracy 0.098 0.6843 0.6846 0.8967 0.8918 0.8653 0.8992 0.9137 0.9155
Optimality Gap 2.0245 1.9230 1.9228 1.4710e-01 1.5514e-01 2.1553e-01 1.2826e-01 5.0790e-02 3.7756e-02
S&PI is the abbreviation of settings and performance indices.
@@
Gaussian attacks: It is shown in Fig. (3a) that an $m=40$ multi-agent CPS consists of ${\left|\mathcal{R}\right|}=32$ reliable agents (yellow nodes) and $\left|\mathcal{B}\right|=8$ Byzantine agents (red nodes), where each Byzantine agent $o$, $o\in\mathcal{B}$, sends a message generated by a Gaussian distribution with mean $\sum\nolimits_{j\in{\mathcal{R}_{i}}}{{w_{ij}}{x_{j,k}}}/\sum\nolimits_{j\in{\mathcal{R}_{i}}}{{w_{ij}}}$ and standard deviation 30, to its reliable neighbor $i$, $i\in{\mathcal{R}_{o}}$ at each iteration. This attack serves as a Gaussian noise, which can easily inflict fluctuation on the status of reliable agents and deviate the states from their true values. Even though the testing accuracy index can still fluctuate around 0.6, we can see from Fig. (3b) and Table III that NIDS and PMGT-SAGA/PMGT-LSVRG show divergence from the optimal solution under Gaussian attacks. This is the gap between theoretical results and experimental performance. It is shown by Figs. (3b)-(3d) that Prox-DBRO-SAGA and Prox-DBRO-LSVRG can still achieve a smaller optimality gap and higher testing accuracy in the same epochs, alternatively faster convergence, than the other tested algorithms. Moreover, one can clearly see from Table III that Prox-DBRO-SAGA takes the superiority on all three performance indices (consensus errors, testing accuracy, and optimality gaps) at 150 epochs, while Prox-DBRO-LSVRG ranks the second leading position on these three performance indices.
@@
(b) Optimality gaps over epochs.
@@
(d) Consensus errors over epochs.
Figure 3: Performance of all tested algorithms under Gaussian attacks.
S&PI Algorithms NIDS PMGT-LSVRG PMGT-SAGA Prox-BRIDGE-T Prox-BRIDGE-M Prox-BRIDGE-K Prox-GeoMed Prox-DBRO-LSVRG Prox-DBRO-SAGA
Step-Size [0.3, 0.35] 0.3 0.3 0.4 0.3 0.3 0.4 0.02 0.0025
Triggered Probability N/A $m/Q$ N/A N/A N/A N/A N/A $\left[{m/Q/4,m/Q/2}\right]$ N/A
Penalty Parameter N/A N/A N/A N/A N/A N/A N/A [0.05, 0.055] [0.005, 0.0055]
Consensus Error 7.7540e+04 3.2041e+04 3.1770e+04 1.7037e-03 1.1220e-03 1.9835e-03 0.9002 0.0007 6.71366e-08
Testing Accuracy 8.1060e-01 4.9850e-01 5.7650e-01 8.9920e-01 8.9540e-01 8.5220e-01 0.0017 0.9182 9.1970
Optimality Gap Inf 3.9609e+01 3.1770e+04 1.2136e-01 1.4870e-01 2.4074e-01 1.1507e-01 4.0442e-02 2.4398e-02
S&PI is the abbreviation of settings and performance indices.
@@
(b) Optimality gaps over epochs.
@@
(d) Consensus errors over epochs.
Figure 4: Performance of all tested algorithms under same-value attacks.
S&PI Algorithms NIDS PMGT-LSVRG PMGT-SAGA Prox-BRIDGE-T Prox-BRIDGE-M Prox-BRIDGE-K Prox-GeoMed Prox-DBRO-LSVRG Prox-DBRO-SAGA
Step-Size [0.5, 0.55] 0.5 0.5 0.2 0.3 0.4 0.4 $0.35/\sqrt{k+1}$ $0.11/\sqrt{k+1}$
Triggered Probability N/A $m/Q$ N/A N/A N/A N/A N/A $\left[{m/Q/8,m/Q/4}\right]$ N/A
Penalty Parameter N/A N/A N/A N/A N/A N/A N/A [0.0005, 0.00055] [0.0005, 0.00055]
Consensus Error 0.0417 1121561.6743 1121561.6458 0.0514 0.00174 190905836.8303 0.0027 0.0006 0.0008
Testing Accuracy 0.098 0.8019 0.8162 0.8023 0.8942 0.8649 0.9006 0.9065 0.9055
Optimality Gap 6.5336e+04 2.2774e+04 2.2774e+04 Inf 1.7867e-01 4.2663e+01 1.6547e-01 1.6316e-01 1.6534e-01
S&PI is the abbreviation of settings and performance indices.

### v12

@@ IV-B Main Results
(Bounded subgradients). The subgradient $\partial_{\tilde{x}}g\left({\tilde{x}}\right)$ at any point ${\tilde{x}}\in\mathbb{R}^{n}$ is $\hat{G}$-bounded, i.e., $\left\|{{\partial_{\tilde{x}}}g\left({\tilde{x}}\right)}\right\|_{2}^{2}\leq\hat{G}$, where $\hat{G}$ can be an arbitrarily large but finite positive constant.
To proceed, we define $\theta>4/\gamma$, $P_{1}^{d}:=16{\left|\mathcal{R}\right|}\hat{G}{\rm{+}}16n{\phi^{2}}\sum\nolimits_{i\in{\mathcal{R}}}{{{\left|{{{\mathcal{R}}_{i}}}\right|}^{2}}}{\rm{+4}}n{\phi^{2}}\sum\nolimits_{i\in{\mathcal{R}}}{{{\left|{{{\cal B}_{i}}}\right|}^{2}}}$, $\Xi:=\max\left\{{\frac{{{\theta^{2}}P_{1}^{d}}}{{\gamma\theta-1}},\left({\xi-\frac{\gamma}{4}\theta}\right){\left\|{{x_{0}}-{x^{*}}}\right\|_{2}^{2}}+\frac{{{\theta^{2}}}}{\xi}P_{1}^{d}+\theta{P_{2}}-\xi E}\right\}$, and $\xi:={\kappa_{q}}\left({64{{\left({1+{\kappa_{f}}}\right)}^{2}}+{q_{\min}}}\right)\mu\theta$.
@@
The convergence results established in Theorems 2-3 assert that the proposed algorithms achieve linear convergence at the expense of a larger convergence error than that of in the sub-linear convergence case. We note that a smaller constant step-size in the linear convergence case may simultaneously lead to a smaller convergence error and a slower convergence rate according to Theorem 2. It is also clear from Theorem 3 that the convergence error of Prox-DBRO-SAGA and Prox-DBRO-LSVRG for the sub-linear convergence case is determined by the number of Byzantine agents. That is to say, the sub-linear exact convergence of Prox-DBRO-SAGA and Prox-DBRO-LSVRG can be recovered, when the number of Byzantine agents equals to zero. Therefore, the theoretical results derived in Theorems 2-3 demonstrate a trade-off between the convergence error and convergence rate, which has been revealed by the theoretical results.
Remark 7
This paper does not make any assumption or restrictions on the number/proportion of Byzantine agents in the network and only assumes a connected network among all reliable agents (see Assumption 2). However, this does not imply that the number of Byzantine agents can be unbounded since an increase in the number of Byzantine agents lead to larger convergence errors in both cases and an unbounded number of Byzantine agents causes eventually divergence of the proposed algorithms according to Eqs. (13) and (14) in Theorems 2-3. According to Theorems 2-3, the resilience of Algorithms 2-3 is characterized by the consensus and controllable convergence errors of all reliable agents.
The sequel corollary verifies the statement that there is no restriction on the number or proportion of Byzantine agents under appropriate assumptions.
Corollary 1
(Guaranteed convergence for a high proportion of Byzantine agents) Suppose that Assumptions 1-2 hold. Under the conditions of Theorem 2 (resp., Theorem 3), if we set the number of Byzantine agents by $\left|{\cal B}\right|={r_{a}}\left|{\cal R}\right|$ with any bounded constant $0<r_{a}<\infty$ that is only need to be properly chosen to avoid floating point issues, then the sequence ${\left\{{{x_{k}}}\right\}_{k\geq 0}}$ generated by Algorithms 2-3 with the constant (resp., decaying) step-size, still converges to an error ball around the optimal solution to the original problem (1) at a linear (resp., sub-linear) rate of ${\left({1-\mathcal{O}\left({\gamma\alpha}\right)}\right)^{k}}$ (resp., $\mathcal{O}\left({1/k}\right)$), $\forall k\geq 1$, where the radius of the error ball is bounded.
Proof 8
We begin by showing that the convergence error is bounded for any $0<r_{a}<\infty$ in the linear convergence case. By relaxing the upper bound in the right-hand-side (RHS) of (13), we have
$\mathbb{E}\left[{\left\|{{x_{k}}-{1_{\left|{\cal R}\right|}}\otimes{{\tilde{x}}^{*}}}\right\|_{2}^{2}}\right]$ (15)
$\leq$ ${\left({1-\frac{\gamma}{4}\alpha}\right)^{k}}{U_{0}}+{4n\left({4\alpha\left({\frac{{4+r_{a}^{2}}}{\gamma}}\right)+{{\left({\frac{{{r_{a}}}}{\gamma}}\right)}^{2}}}\right){{\phi}^{2}}{{\left|{\cal R}\right|}^{3}}}.$
where the first term in the RHS of (15) decays linearly over the iteration $k$, $\forall k\geq 1$, as $10<1-\gamma\alpha/4<$ 1 according to Theorem 2 and $U_{0}$ is a constant owing to the initialization rule of the proposed algorithms. Moreover, since both $n$, $\gamma$, $\phi$, and the step-size $\alpha$ are (bounded) constants once the objective function and network are determined, we know that the second term, i.e., the error term, in the RHS of (15) can be bounded by a finite value when $0<r_{a}<\infty$. Therefore, the radius of the convergence error ball is bounded. We omit the proof for the sub-linear case since it follows a same technical line as in the linear case.
Remark 8
Corollary 1 asserts that the radius of the error ball is bounded without any restriction on the number or proportion of Byzantine agents under some appropriate conditions. A specific example is when the proportion of Byzantine agents accounts for 50%, i.e., $r_{a}=1$, the radius of the convergence error ball is $4n\left({20\gamma\alpha+1}\right){{\phi}^{2}}{\left|{\cal R}\right|^{3}}/{\gamma^{2}}$ according to (15), which is bounded since all the parameters are constants once the objective function and network are determined. However, this does not imply that the number of Byzantine agents can be arbitrarily large since an increase in the number of Byzantine agents lead to larger convergence errors in both cases and an unbounded number, i.e., ${r_{a}}\to\infty$, of Byzantine agents causes eventually divergence of the proposed algorithms according to Theorems 2-3
V Numerical Experiments
@@
In this section, we perform a case study on decentralized soft-max regression with sparsity to verify the theoretical results and show the convergence performance of the proposed algorithms, where four kinds of Byzantine attacks (zero-sum attacks, Gaussian attacks, same-value attacks, and sign-flipping attacks) are considered. The communication networks are randomly generated by the Erdős-Rényi method, where Byzantine agents are also selected in a random way. Most existing literature adopts only testing accuracy and the consensus error to validate the convergence performance of their proposed algorithms. However, neither higher testing accuracy nor a smaller consensus error can comprehensively reflect the convergence of the tested algorithms. This is because these two metrics fail to precisely measure the iterative distance between the optimized function value and the optimal value of optimization problems, which, however, serves as a primary goal of theoretical analysis. Therefore, there is a gap between the theoretical result and practical performance. To bridge the gap, we introduce the (averaged) optimal gap in the form of function values, i.e., $\left({1/\left|\mathcal{R}\right|}\right)\sum\nolimits_{i\in\mathcal{R}}{\left({{f_{i}}\left({{x_{i,k}}}\right)+g\left({{x_{i,k}}}\right)-\left({{f_{i}}\left({{{\tilde{x}}^{*}}}\right)+g\left({{{\tilde{x}}^{*}}}\right)}\right)}\right)}$, as the third metric, which could precisely captures transient behaviors (convergence or divergence) of algorithms when training a machine- or deep-learning model. A network of $m$ agents consisting of ${\left|\mathcal{R}\right|}$ reliable agents and $\left|\mathcal{B}\right|$ Byzantine agents, minimize a regularized soft-max regression problem for a multi-class classification task via specifying the problem formulation (1) as ${f_{i}}\left({\tilde{x}}\right):=-\left({1/{q_{i}}}\right)\sum\nolimits_{j=1}^{{q_{i}}}{\sum\nolimits_{l=0}^{\tilde{C}-1}{{\mathcal{I}_{\left({{\tilde{b}_{ij}}=l}\right)}}}}\ln\left({{e^{\left[{\tilde{x}}\right]_{l}^{\top}{\tilde{c}_{ij}}}}/\sum\nolimits_{t=0}^{\tilde{C}-1}{{e^{\left[{\tilde{x}}\right]_{t}^{\top}{\tilde{c}_{ij}}}}}}\right)+\left({{\beta_{1}}/2}\right)\left\|{\tilde{x}}\right\|_{2}^{2}$ and ${g}\left({\tilde{x}}\right):={\beta_{2}}{\left\|{\tilde{x}}\right\|_{1}}$, where $\tilde{x}\in{\mathbb{R}^{n}}$ with $n=\tilde{C}\tilde{n}$ is the model parameter, ${\tilde{C}}$ represents the number of sample classes, ${\left[{\tilde{x}}\right]_{j}}$ denotes a vector that contains $\left(j\tilde{n}\right)$-th to $\left(\left({j+1}\right)\tilde{n}-1\right)$-th elements of $\tilde{x}$, ${{{\tilde{b}_{ij}}}}$ and ${{\tilde{c}_{ij}}}$ are the $j$-th label and image allocated to agent $i$, respectively, ${{\mathcal{I}_{\left({{{\tilde{b}_{ij}}}=l}\right)}}}$ is the indicator function with ${\mathcal{I}_{\left({{{\tilde{b}_{ij}}}=l}\right)}}=1$ if ${{{\tilde{b}_{ij}}}=l}$ and ${\mathcal{I}_{\left({{{\tilde{b}_{ij}}}=l}\right)}}=0$ otherwise, ${\beta_{1}}$ and ${\beta_{2}}$ are positive parameters of regularized terms for avoiding over-fitting and obtaining a sparse solution, respectively. We denote the total number of training samples by $N$ and the regularized parameters are set as ${\beta_{1}}={\beta_{2}}=1/N$ in the following numerical experiments.
@@
Since the algorithmic framework BRIDGE [23] and decentralized algorithm (denoted by Peng) [27] are only available to handling a class of smooth single-objective optimization problems, we equip them with the proximal-gradient mapping method [39, 40, 41, 38, 12] to obtain Prox-BRIDGE-T, Prox-BRIDGE-M, Prox-BRIDGE-K, and Prox-Peng for the nonsmooth composite finite-sum optimization problem, which is also applied to GeoMed [53] to get Prox-GeoMed. The initial state of decision variables of all tested algorithms are the same and generated from a standard normal distribution. Note that the parameters of all tested algorithms are optimized manually to obtain their best performance, and the parameters associated with the problem model keep the same to ensure fairness. A total number of $Q=60000$ training samples from the MNIST [54] data set are evenly allocated to each agent (including both reliable agents and Byzantine agents in the network) to train the discriminator, while the rest $10000$ samples are used for testing. Fig. 1 presents 100 samples randomly selected from the data set. Recall the theoretical results regarding the decaying and constant step-sizes and penalty parameter such that the experimental setting gives the following feasible selection ranges: $\alpha\in\left({0,0.1385}\right]$, ${\alpha_{k}}\in\left({0,1/\left({k+14.2733}\right)}\right]$, and $\phi\geq 0.0003$.
@@
(b) Optimal gap over epochs.
@@
(d) Consensus error over epochs.
Figure 2: Performance of all tested algorithms under zero-sum attacks.
@@
PS&PI Algorithms NIDS PMGT-LSVRG PMGT-SAGA Prox-BRIDGE-T Prox-BRIDGE-M Prox-BRIDGE-K Prox-GeoMed Prox-Peng Prox-DBRO-LSVRG Prox-DBRO-SAGA
Step-size [0.01, 0.015] 0.001 0.01 0.35 0.3 0.35 0.35 0.5 0.05 0.005
Triggering probability N/A $m/Q$ N/A N/A N/A N/A N/A N/A $\left[{m/Q/2,m/Q}\right]$ N/A
Penalty parameter N/A N/A N/A N/A N/A N/A N/A 0.1 [0.2, 0.25] [0.2, 0.25]
Consensus error 0 1.4163e-11 1.4159e-09 1.2492e-03 1.2116e-03 1.9340e-03 3.8010e-03 92.9523 4.4680 4.2898e-02
Testing accuracy 0.098 0.6812 0.6812 0.8965 0.8918 0.8653 0.8999 0.8789 0.9137 0.9155
Optimal gap 2.0245 1.9230 1.9228 1.4940e-01 1.5751e-01 2.1701e-01 1.2435e-01 2.7077e-01 5.0790e-02 3.7756e-02
PS&PI is the abbreviation of parameter settings and performance metrics.
Zero-sum attacks: As depicted in Fig. 2: LABEL:sub@Fig-1-1, an $m=30$ network consists of ${\left|\mathcal{R}\right|}=25$ reliable agents (yellow nodes) and $\left|\mathcal{B}\right|=5$ Byzantine agents (red nodes), where each Byzantine agent $b$, $b\in\mathcal{B}$, sends a well-designed malicious message ${z_{ib,k}}=-{\sum\nolimits_{j\in{\mathcal{R}_{i}}}{{w_{ij}}{x_{j,k}}}}/\left|{{\mathcal{B}_{i}}}\right|/{w_{bi}}$ to its reliable neighbor $i$, $i\in{\mathcal{R}_{b}}$, to drive the states of the reliable agent ${x_{i,k}}=0_{n}$ at each iteration. For NIDS [40], the algorithm parameter is fixed as $c=1/\left({2{{\max}_{i\in\mathcal{R}}}\left\{{{\alpha_{i}}}\right\}}\right)$ for $\tilde{W}={I_{m}}-c{D_{\alpha}}\left({{I_{m}}-W}\right)$, where $\tilde{W}$, $W$, and $D_{\alpha}$ are the modified mixing matrix, weight matrix, and uncoordinated step-size, respectively. This means if $c$ is sufficiently small, NIDS runs without any communication happening among all agents (both reliable agents and Byzantine agents) in the network. For PMGT-SAGA/PMGT-LSVRG [37], we hand-tune the parameter associated with multi-step communications to obtain the best performance. It is clear from Fig. 2: LABEL:sub@Fig-1-2-LABEL:sub@Fig-1-4 and Table II that the proposed algorithms achieve a smaller optimal gap and higher testing accuracy than the other tested algorithms in a same amount of computational costs (epochs). This demonstrates that the proposed algorithms approximate faster to the optimal solution than the other tested algorithms. Notably, zero-sum attacks launched by Byzantine agents aim to drive the states of all reliable agents to zero at each iteration. Therefore, a much smaller consensus error of NIDS and PMGT-SAGA/PMGT-LSVRG than the other Byzantine-resilient decentralized algorithms, may indicate that they are less resilient or more susceptible to the zero-sum attacks. This can be testified by their bigger optimal gaps and lower testing accuracy shown in Fig. 2: LABEL:sub@Fig-1-2-LABEL:sub@Fig-1-3 and Table II.
@@
(b) Optimal gap over epochs.
@@
(d) Consensus error over epochs.
Figure 3: Performance comparison among all tested algorithms under Gaussian attacks.
Gaussian attacks: It is shown in Fig. 3: LABEL:sub@Fig-2-1 that an $m=40$ network consists of ${\left|\mathcal{R}\right|}=32$ reliable agents (yellow nodes) and $\left|\mathcal{B}\right|=8$ Byzantine agents (red nodes), where each Byzantine agent $b$, $b\in\mathcal{B}$, sends a message subject to a Gaussian distribution with mean $\sum\nolimits_{j\in{\mathcal{R}_{i}}}{{w_{ij}}{x_{j,k}}}/\sum\nolimits_{j\in{\mathcal{R}_{i}}}{{w_{ij}}}$ and standard
@@
PS&PI Algorithms NIDS PMGT-LSVRG PMGT-SAGA Prox-BRIDGE-T Prox-BRIDGE-M Prox-BRIDGE-K Prox-GeoMed Prox-Peng Prox-DBRO-LSVRG Prox-DBRO-SAGA
Step-size [0.3, 0.35] 0.3 0.3 0.4 0.3 0.3 0.4 0.05 0.0015 0.0025
Triggering probability N/A $m/Q$ N/A N/A N/A N/A N/A N/A $\left[{m/Q/4,m/Q/2}\right]$ N/A
Penalty parameter N/A N/A N/A N/A N/A N/A N/A 0.001 [0.001, 0.0015] [0.001, 0.0015]
Consensus error 3.0396e+04 3.2041e+04 3.1770e+04 1.8532e-03 1.0682e-03 1.9835e-03 1.6716e-03 8.3993e-02 1.1938e-03 5.1252e-06
Testing accuracy 0.0742 0.4985 0.5765 0.9001 0.8952 0.8522 0.9002 0.8856 0.9165 0.9197
Optimal gap Inf 4.2709e+01 3.7046e+01 1.2136e-01 1.4870e-01 2.4074e-01 1.1507e-01 1.4480e-01 3.2053e-02 5.7316e-03
PS&PI is the abbreviation of parameter settings and performance metrics.
deviation 30, to its reliable neighbor $i$, $i\in{\mathcal{R}_{b}}$ at each iteration. This attack serves as a Gaussian noise, which can easily inflict fluctuation on the state of reliable agents and deviate the states from their true values. Even though the testing accuracy index can still fluctuate around 0.6, we can see from Fig. 3: LABEL:sub@Fig-2-2 and Table III that NIDS and PMGT-SAGA/PMGT-LSVRG show divergence from the optimal solution under Gaussian attacks. It is shown by Figs. 3: LABEL:sub@Fig-2-2-LABEL:sub@Fig-2-4 that the proposed algorithms can still achieve a smaller optimal gap and higher testing accuracy in the same epochs, alternatively faster convergence, than the other tested algorithms. Moreover, one can clearly see from Table III that Prox-DBRO-SAGA takes the superiority on all three performance metrics (optimal gap, testing accuracy, and consensus error) at 150 epochs, while Prox-DBRO-LSVRG ranks second on these three performance metrics.
@@
(b) Optimal gap over epochs.
@@
(d) Consensus error over epochs.
Figure 4: Performance comparison among all tested algorithms under same-value attacks.
@@
PS&PI Algorithms NIDS PMGT-LSVRG PMGT-SAGA Prox-BRIDGE-T Prox-BRIDGE-M Prox-BRIDGE-K Prox-GeoMed Prox-Peng Prox-DBRO-LSVRG Prox-DBRO-SAGA
Step-size [0.5, 0.55] 0.5 0.5 0.35 0.3 0.4 0.4 $0.97/\left({k+25}\right)$ $0.91/\left({k+21}\right)$ $0.74/\left({k+35}\right)$
Triggering probability N/A $m/Q$ N/A N/A N/A N/A N/A N/A $\left[{m/Q/8,m/Q/4}\right]$ N/A
Penalty parameter N/A N/A N/A N/A N/A N/A N/A 0.0005 [0.0004, 0.00045] [0.0005, 0.00055]
Consensus error 4.1710e-02 1.1216e+06 1.1216e+06 2.0447e-03 1.7397e-03 1.9091e+08 2.6619e-03 6.0639e-01 1.4196e-03 7.0507e-04
Testing accuracy 0.098 0.8227 0.8162 0.8972 0.8942 0.8649 0.9006 0.8915 0.9072 0.9067
Optimal gap 6.5336e+04 2.2783e+04 2.2783e+04 1.6548e-01 1.6724e-01 1.6336e+03 1.5303e-01 1.9619e-01 1.4623e-01 1.5938e-01
PS&PI is the abbreviation of parameter settings and performance metrics.